# Experimentally Testing AI-Powered Content Transformations on Student Learning

### v1

@@ 5.3. Learning Experience
$p<\textbf{0.001}^{***}$
$r=0.49$
@@
$p=\textbf{0.012}^{*}$
$r=0.29$
I would like to use today’s educational tool to
support my learning needs in the future.
4.43 (0.73) 3.83 (1.05)
$Z=2.42$
$p=\textbf{0.008}^{**}$
$r=0.31$
I would recommend this educational tool to
other students to support their learning needs.
4.53 (0.78) 3.97 (0.89)
$Z=2.88$
$p=\textbf{0.002}^{**}$
$r=0.37$
$Z=2.44$
$p=\textbf{0.007}^{**}$
$r=0.31$
@@
$p=\textbf{0.027}^{*}$
$r=0.25$
The educational tool I used today would make me more effective at learning compared to other educational tools I currently use at home or in school. 4.17 (0.79) 3.07 (0.98)
$Z=4.11$
$p<\textbf{0.001}^{***}$
$r=0.53$
6. Qualitative Findings
@@
6.1. RQ1 and RQ2: Learning Outcomes
6.1.1. Dynamic representations of content can effectively help students learn.
One of the unique attributes of Learn Your Way is that it provides multiple interactive transformations in one learning experience. This approach is based on established learning principles, such as Mayer and Moreno’s “Cognitive Theory of Multimedia Learning” (Mayer and Moreno, 1998) and Paivio’s “Mental representations: A dual coding approach,” (Paivio, 1990) which posit that encoding information in multiple formats can make learning more effective by engaging multiple senses and cognitive channels, and by simplifying complex ideas. Students in the present study expressed value in the ability to access multiple learning modalities. These were perceived as useful for clarifying concepts and improving comprehension and retention of the learning material:
“Reading it multiple times but in different ways made it stick in my mind. In the assessment, some topics I remembered from the slide, some I remembered from reading. So, reading it multiple times but in different ways made it stick in my mind.” - P86, Urban, Learn Your Way
“If I didn’t feel I was getting good information from one [modality], I could go to the next one… Immersive Text and the slides are really good for a study tool especially. Mindmap helped me put everything together and really see where everything fits in here… I think because it’s the same information, you just kind of can play around with it and go to different things like Slides and stuff. I think that, in itself, helped me to remember [the material] and see what’s really going on.” - P20, Suburban, Learn Your Way
Conversely, our qualitative feedback revealed that participants from the Digital Reader condition desired the ability to view content in multiple formats. Despite being unaware of the Learn Your Way condition, several participants in the Digital Reader condition often described features in Learn Your Way when asked what their ideal learning tool would be:
“I’m more of a visual learner, so if I see something visual, I better understand it than words or text.” - P09, Urban, Digital Reader Condition
@@ 6.1. RQ1 and RQ2: Learning Outcomes
6.1.2. Summaries and smaller chunks of text via Immersive Text in Learn Your Way improved students’ comprehension of the content
@@
“I do feel the way that they structured it - having only a few paragraphs or split up text definitely helps my brain not get overwhelmed by the whole thing. So not just putting a whole chunk of text that is relating to the whole topic and then just being like, ‘okay here.’” - P40, Suburban, Learn Your Way
“Like anything in my life, if I see something I have to tackle…and it’s one big thing - I struggle breaking it up into smaller parts. If it’s reading a really long article for a class or cleaning my room, I look at it and I’m like, ‘whoa, that’s a lot of stuff. I can’t do it.’ If it’s broken up for me, I feel less overwhelmed by it.” - P13, Suburban, Learn Your Way
Some students in the Digital Reader condition however, felt overwhelmed by having only the chapter’s content, which made it harder to focus and retain information. We also observed that some students in this condition attempted to create summaries and break the content up on their own by using the annotation tools in the digital reader.
“Because even though the paragraphs aren’t that big, there’s still enough paragraphs where you’ll forget your information” - P54, Rural, Digital Reader Condition
“Maybe if it stated the key points a little better because only about half of the article has like an actual head headliner, you know, forgot what they’re called…so yeah, I mean maybe just like breaking down the big chunks of text a little better.” -P94, Suburban, Digital Reader Condition
@@
6.1.3. The value of assessments via QuizMe was a key driver in performance outcomes
@@
One of the helpful ways students effectively learned the material in Learn Your Way was through the QuizMe feature. All participants in Learn Your Way acknowledged the role of quizzes and feedback on their performance. This is consistent with previous research that demonstrates that quizzes can significantly enhance learning outcomes by offering additional practice exercises in areas where a student is struggling (Chang and Chien, 2024). This approach helps learners attain mastery at their own pace and speeds up the learning process (Al-Badi et al., 2022; Kulik and Fletcher, 2016). Quizzes were a favored way to test comprehension of the material and feedback instilled confidence in learning by increasing perceptions of knowledge retention for students in the Learn Your Way condition.
@@
“The quizzes for sure helped me get the right answers. It helped me even if I needed help, it gave a hint if I didn’t understand it. I feel like it also made it really easy to explain or understand.” - P20, Urban, Learn Your Way
@@
“I liked how it gave me feedback and told me why I got it right versus just bringing me to the next question right away. It was helpful and made me realize how my thought process was and how I got to the answer.” - P11, Suburban, Learn Your Way
@@
In contrast, students in the Digital Reader condition reported missing the ability to check their understanding and receive feedback through guided activities and tools such as quizzes, flashcards and look up tools:
@@
“I would probably make a tool where everything is together, like the test taking with the text so I can select what I want to learn more about in the text and put it in a test where I’m able to answer the questions if I know it goes away, if I don’t I have to keep studying it and learn more.” - P12, Suburban, Digital Reader Condition
@@
“Like the ideal learning tool, like if I were to learn this article a different way, I mean, yeah, like I said, I mean flashcards I find useful. I definitely think once you read it, once you like read a certain amount of it or whatever, like certain checkpoints I guess it could quiz you on what you’ve done so far to review to make sure that you got all the key points.” - P94, Suburban, Digital Reader Condition
@@
Feedback from participants aligns with educational theories about active engagement in learning and emphasizes the need to provide active checkpoints in learning experiences (Moreno and Mayer, 2000). This approach allows students to leverage metacognitive strategies to reflect on their understanding of the learning material and increase the likelihood of retaining information (Zimmerman and Moylan, 2009).
6.2. RQ3: Learning Experiences
@@
6.2.1. The variety of capabilities and self-directed nature of Learn Your Way played a critical role in creating a positive learning experience
@@
Access to multiple transformations and the freedom to choose which ones to learn with resulted in an engaging learning experience for students in the Learn Your Way condition. This is also consistent with research that advocates for learners to play a central role in their own learning journeys (Deci and Ryan, 2000; Zimmerman and Moylan, 2009; Zimmerman, 2000).
@@
“I personally like going in my own structured way… It’s just a learning preference… I like it more of me leading it because I feel more in control of my learning…it allows me to go my own pace and let me learn how I want to learn.” - P08, Urban, Learn Your Way
@@
“I like how many options it has for different ways of taking in the information because I definitely like learning it one way and then being able to pick another way to review…” - P13, Suburban, Learn Your Way
@@
Conversely, students in the Digital Reader condition who did not have access to transformations, noted the challenges of this learning experience and desired more interactions.
@@
“I guess I liked how it was kind of just all there and I could read it. What I didn’t like…It was just like kind of boring reading it…I was just like sitting there trying to memorize it. Reading it…I usually learn best like videos and stuff. I can’t like read stuff and store in memory…I guess I’m like more of like a visual learner.” - P19, Suburban, Digital Reader Condition
@@
6.2.2. Auditory transformations were perceived as useful for combining learning with other activities
@@
“I do listen to like a lot of podcasts in general, but I always like to have like, stuff playing while I’m doing stuff. So like, like I was saying before, like, oh, if you’re like cooking or cleaning your room or just doing anything really like, I don’t know, like even like hobbies, like you’re like drawing or something. Like just having, I always like to have something playing in the background. So like when I have the option to have that for like school stuff, I just like have, I feel it feels more productive. Like I can get it done while I’m doing other stuff, while like retaining the information as well” - P86, Urban, Learn Your Way
@@
“I mean it’s the same with the Video I guess. It’s like, like people talking rather than somebody explaining they’re, I mean, in my mind pretty similar in use case. Like if I’m driving or like, you know, cleaning or doing something where I can’t like read and I can just throw this on and learn about it. That would be the use case” - P01, Urban, Learn Your Way
@@
The Digital Reader on the other hand, required more visual focus to process the content in the textbook chapter. Without access to text to speech tools or additional support, multitasking with the Digital Reader would therefore be challenging. It is worth noting however that there are open questions about the impact of multitasking on the comprehension and recall of information and which tasks are suitable for this (Coens et al., 2011). It is therefore crucial to consider approaches to support metacognitive strategies (e.g. reflection) with auditory modalities to ensure effective learning.
7. Discussion
@@
Our results are in support of our hypotheses, which demonstrate that Learn Your Way can lead to both improved learning outcomes and experiences. AI-powered content transformation tools can be a powerful driver of pedagogical innovation. In particular, our experiment reveals that these tools are not merely supplementary aids but can play a critical role in reshaping how students interact with and comprehend academic content. While many studies exploring the impact of AI tools on educational outcomes have been mixed (e.g.,(Do et al., 2025; Sun et al., 2024)), our research demonstrates that Learn Your Way can enhance students’ immediate and long-term recall of information and create a more effective learning environment compared to a traditional digital reader.
@@
The experimental results suggest that this can be attributed to several factors. First, learning is a complex and multifaceted process and students exhibit diverse preferences in how they process information and acquire knowledge, including visual, kinesthetic, and reading/writing approaches (MCcall, 2024). Traditional instructional methods often fail to address these unique needs at scale, leading to suboptimal educational experiences (Villegas-Ch et al., 2024). Instead, engaging with information through multiple representational formats, such as slide decks, audio lessons, or visual mind-maps, facilitates the construction of a more robust and comprehensive mental model of the content (Ainsworth, 1999). This is because diverse representations can clarify complex ideas and highlight distinct facets of a concept. Learn Your Way addresses this need by offering students the ability to select their preferred learning modalities, enabling them to explore concepts from a range of engaging and diverse formats. Second, simplified and multimodal presentations can reinforce learning, leading to greater retention when information is seen, heard, and actively engaged with (Clark and Paivio, 1991). By breaking down content into manageable chunks and allowing students to personalize the presentation of the content in dynamic formats based on their learning preferences, we believe students forged better mental connections between different representations, resulting in better comprehension of the content compared to students who engaged with the digital reader.
@@
Drawing from established literature on the benefits of adaptive quizzing and personalized feedback (Frank, 2024), this study’s quantitative and qualitative findings underscore their critical role in the learning experience. All students who used Learn Your Way experienced quizzing and feedback. Participants’ qualitative feedback further highlighted the value of these capabilities, with many emphasizing how quizzes aided in testing comprehension and retention, while the associated feedback was instrumental in helping them understand the rationale behind their answers.
@@
Given the power of content transformation capabilities on student learning, one might wonder why there was no direct correlation between engagement with usage of more content transformation modalities and better learning outcomes among students who used Learn Your Way. This analysis was exploratory and the pattern of data is in the direction we expected (i.e., more use of Learn Your Way modalities was correlated with better learning outcomes), but not statistically significant. We expect this may have occurred for two reasons. First, this analysis required us to segment our Learn Your Way students into two different groups, which significantly reduced our sample size and therefore power to detect statistical significance. Second, our qualitative feedback suggests that the usage of additional modalities, such as Slides and Video were often treated as supplementary resources for content review, following primary engagement with the Immersive Text and QuizMe features.
@@
“I feel like (Slides) would be good for review after I already read the Immersive Text if I wanted to review the main points.” - P72, Rural
@@
Our findings therefore suggest two potential explanations for the observed phenomena. First, a larger sample size could provide a more robust basis for investigating whether increased use of content transformation capabilities positively impacts learning outcomes. Second, it may be the case that a greater quantity of engagement with content transformation features does not, in itself, lead to improved learning outcomes. Rather, our results corroborate previous research demonstrating that enhanced learning is contingent upon the strategic alignment of an individual’s learning preferences with the specific demands of the task (Cordova and Lepper, 1996; Schneider et al., 2018). In other words, a critical consideration in implementing content transformation capabilities is ensuring student agency. While these tools offer significant value, their effectiveness is maximized when learners have the autonomy to select the specific transformations that align with their personal learning goals and contextual needs. Therefore, the effectiveness of content transformation tools may not be predicated on the sheer number of modalities utilized but rather on the relevance and quality of student interactions with those modalities within a given learning context. Based on the qualitative feedback from our study where participants emphasized the value of different content transformations when learning the chapter, we posit the former explanation is more likely. However, future research would be helpful in order to clarify causal mechanisms.
@@
Our results also demonstrate that students who engaged with Learn Your Way reported a more positive learning experience compared to those who used a traditional digital reader. The present study’s findings are supported by a body of literature suggesting that transforming educational content into multiple representations can make the learning process more enjoyable, engaging, and motivating (Ainsworth, 1999). This is particularly powerful when students are able to select representations that best match their individual learning preferences (Surjono, 2015; Popescu, 2010). Learn Your Way, which is capable of dynamically transforming content, represents a critical evolution from standard digital textbooks by enabling more tailored and engaging educational experiences for students.
@@
7.1. Limitations and Future Work
While our findings provide evidence that AI-powered content transformation tools can positively impact learning outcomes and experiences, our study is not without limitations. First, our study took place in an experimental lab setting. Future research should explore best practices for how and when to utilize these content transformation tools within a classroom environment, and collect longitudinal outcomes on learning and retention. Second, our study does not isolate which, if any, specific capabilities present within Learn Your Way lead to improved outcomes. Rather, we demonstrate positive results from the experience overall. Further research is needed to explore how specific content transformations may, or may not, interact with the design of the user experience to impact learning outcomes. Finally, our research covered one subject area with high school aged participants. More research is needed to explore how content transformations impact learning across a variety of school subjects and a wider range of student ages.
8. Conclusion

### v2

@@ 6.3 Learning Experience
$p_{adj}<.001^{***}$
$r=.49$
The educational tool I used today would make me more effective at learning compared to other educational tools I currently use at home / school. 4.17 (0.79) 3.07 (0.98)
$Z=4.11$
$p_{adj}<.001^{***}$
$r=.53$
I would recommend this educational tool to
other students to support their learning needs.
4.53 (0.78) 3.97 (0.89)
$Z=2.88$
$p_{adj}=.004^{**}$
$r=.37$
$Z=2.44$
$p_{adj}=.01^{*}$
$r=.31$
I would like to use today’s educational tool to
support my learning needs in the future.
4.43 (0.73) 3.83 (1.05)
$Z=2.42$
$p_{adj}=.01^{*}$
$r=.31$
@@
$p_{adj}=.014^{*}$
$r=.29$
@@
$p_{adj}=.027^{*}$
$r=.25$
In support of our second hypothesis (H2), we observed meaningful differences when we analyzed students’ reported experience with the educational tools. Students in the Learn Your Way condition ($Mdn=4$; $M=3.6$, $SD=0.6$) found the tool to be more useful for learning the educational content compared to Textbook students ($Mdn=3$; $M=3.1$, $SD=0.7$). Specifically, a one-sided Wilcoxon-Mann-Whitney test revealed that Learn Your Way participants rated the educational tool significantly higher on a 4-point Likert scale than Textbook participants, before and after correcting for multiple comparisons with other experience metrics ($Z=3.08$, $p=.001$, $p_{adj}=.003$, $r=.40$).
Furthermore, through a collection of agreement statements about the educational tool, we can see how the Learn Your Way platform positively impacted students’ perceived learning experience as compared to students in the Textbook condition (as shown in Table 7). Specifically, students in the Learn Your Way condition reported higher levels of agreement compared to students in the Textbook condition across all of our experiential measures. Directional Wilcoxon-Mann-Whitney tests were conducted on each statement’s agreement ratings, followed by FDR corrections for multiple comparisons. Table 7 provides a full breakdown of the included statements and mean values by educational tool condition, as well as the associated test statistics with the adjusted p-values reported. The most significant differences revealed between the Learn Your Way platform and the Textbook condition suggest Learn Your Way has the strongest influence on students’ comfort in test-taking, and their perception of the tool’s effectiveness for learning compared to other educational tools already being used at home or in their classroom.
7 Qualitative Findings
@@
7.1 RQ1 and RQ2: Learning Outcomes
7.1.1 Dynamic representations of content can effectively help students learn.
One of the unique attributes of Learn Your Way is that it provides multiple interactive transformations in one learning experience. This approach is based on established learning principles, such as Mayer and Moreno’s “Cognitive Theory of Multimedia Learning” (mayer1998cognitive) and Paivio’s “Mental representations: A dual coding approach,” (Paivio1990) which posit that encoding information in multiple formats can make learning more effective by engaging multiple senses and cognitive channels, and by simplifying complex ideas. Students in the present study expressed value in the ability to access multiple learning modalities. These were perceived as useful for clarifying concepts and improving comprehension and retention of the learning material:
“Reading it multiple times but in different ways made it stick in my mind. In the assessment, some topics I remembered from the slide, some I remembered from reading. So, reading it multiple times but in different ways made it stick in my mind.”
- P86, Urban, Learn Your Way
“If I didn’t feel I was getting good information from one [modality], I could go to the next one… Immersive Text and the slides are really good for a study tool especially. Mindmap helped me put everything together and really see where everything fits in here… I think because it’s the same information, you just kind of can play around with it and go to different things like Slides and stuff. I think that, in itself, helped me to remember [the material] and see what’s really going on.”
- P20, Suburban, Learn Your Way
Conversely, our qualitative feedback revealed that participants from the Textbook condition desired the ability to view content in multiple formats. Despite being unaware of the Learn Your Way condition, several Textbook participants described features already included in Learn Your Way when asked what their ideal learning tool would look like:
“I’m more of a visual learner, so if I see something visual, I better understand it than words or text.”
- P09, Urban, Textbook
@@ 7.1 RQ1 and RQ2: Learning Outcomes
7.1.2 Summaries and smaller chunks of text via Immersive Text in Learn Your Way improved students’ comprehension of the content
@@
“I do feel the way that they structured it - having only a few paragraphs or split up text definitely helps my brain not get overwhelmed by the whole thing. So not just putting a whole chunk of text that is relating to the whole topic and then just being like, ‘okay here.’”
- P40, Suburban, Learn Your Way
“Like anything in my life, if I see something I have to tackle…and it’s one big thing - I struggle breaking it up into smaller parts. If it’s reading a really long article for a class or cleaning my room, I look at it and I’m like, ‘whoa, that’s a lot of stuff. I can’t do it.’ If it’s broken up for me, I feel less overwhelmed by it.”
- P13, Suburban, Learn Your Way
In contrast, some Textbook students felt overwhelmed by having only the chapter’s content, which made it harder to focus and retain information. We also observed that some students in this condition attempted to create summaries and break the content up on their own by using the annotation tools within the digital textbook.
“Because even though the paragraphs aren’t that big, there’s still enough paragraphs where you’ll forget your information.”
- P54, Rural, Textbook
“Maybe if it stated the key points a little better because only about half of the article has like an actual head headliner, you know, forgot what they’re called…so yeah, I mean maybe just like breaking down the big chunks of text a little better.”
-P94, Suburban, Textbook
@@
7.1.3 The value of assessments via QuizMe was a key driver in performance outcomes
@@
One of the helpful ways students effectively learned the material in Learn Your Way was through the QuizMe feature. All participants in Learn Your Way acknowledged the role of quizzes and feedback on their performance. This is consistent with previous research that demonstrates that quizzes can significantly enhance learning outcomes by offering additional practice exercises in areas where a student is struggling (ChangChien2024). This approach helps learners attain mastery at their own pace and speeds up the learning process (al2022perceptions; Kulik2016). Quizzes were a favored way to test comprehension of the material, and feedback instilled confidence in learning by increasing perceptions of knowledge retention for students in the Learn Your Way condition.
@@
“The quizzes for sure helped me get the right answers. It helped me even if I needed help, it gave a hint if I didn’t understand it. I feel like it also made it really easy to explain or understand.”
- P20, Urban, Learn Your Way
@@
“I liked how it gave me feedback and told me why I got it right versus just bringing me to the next question right away. It was helpful and made me realize how my thought process was and how I got to the answer.”
- P11, Suburban, Learn Your Way
@@
In contrast, students in the Textbook condition reported missing the ability to check their understanding and receive feedback through guided activities and tools such as quizzes, flashcards and look up tools:
@@
“I would probably make a tool where everything is together, like the test taking with the text so I can select what I want to learn more about in the text and put it in a test where I’m able to answer the questions if I know it goes away, if I don’t I have to keep studying it and learn more.”
- P12, Suburban, Textbook
@@
“Like the ideal learning tool, like if I were to learn this article a different way, I mean, yeah, like I said, I mean flashcards I find useful. I definitely think once you read it, once you like read a certain amount of it or whatever, like certain checkpoints I guess it could quiz you on what you’ve done so far to review to make sure that you got all the key points.”
- P94, Suburban, Textbook
@@
Feedback from participants aligns with educational theories about active engagement in learning and emphasizes the need to provide active checkpoints in learning experiences (Moreno2000). This approach allows students to leverage metacognitive strategies to reflect on their understanding of the learning material and increase the likelihood of retaining information (Zimmerman2009).
7.2 RQ3: Learning Experiences
@@
7.2.1 The variety of capabilities and self-directed nature of Learn Your Way played a critical role in creating a positive learning experience
@@
Access to multiple transformations and the freedom to choose which ones to learn with resulted in an engaging learning experience for students in the Learn Your Way condition. This is also consistent with research that advocates for learners to play a central role in their own learning journeys (Deci2000; Zimmerman2009; zimmerman2000).
@@
“I personally like going in my own structured way… It’s just a learning preference… I like it more of me leading it because I feel more in control of my learning…it allows me to go my own pace and let me learn how I want to learn.”
- P08, Urban, Learn Your Way
@@
“I like how many options it has for different ways of taking in the information because I definitely like learning it one way and then being able to pick another way to review…”
- P13, Suburban, Learn Your Way
@@
Conversely, students in the Textbook condition who did not have access to transformations, noted the challenges of this learning experience and desired more interactions.
@@
“I guess I liked how it was kind of just all there and I could read it. What I didn’t like…it was just like kind of boring reading it… I was just like sitting there trying to memorize it. Reading it…I usually learn best like videos and stuff. I can’t like read stuff and store in memory… I guess I’m like more of like a visual learner.”
- P19, Suburban, Textbook
@@
7.2.2 Auditory transformations were perceived as useful for combining learning with other activities
@@
“I do listen to like a lot of podcasts in general, but I always like to have like, stuff playing while I’m doing stuff. So like, like I was saying before, like, oh, if you’re like cooking or cleaning your room or just doing anything really like, I don’t know, like even like hobbies, like you’re like drawing or something. Like just having, I always like to have something playing in the background. So like when I have the option to have that for like school stuff, I just like have, I feel it feels more productive. Like I can get it done while I’m doing other stuff, while like retaining the information as well.”
- P86, Urban, Learn Your Way
@@
“I mean it’s the same with the Video I guess. It’s like, like people talking rather than somebody explaining they’re, I mean, in my mind pretty similar in use case. Like if I’m driving or like, you know, cleaning or doing something where I can’t like read and I can just throw this on and learn about it. That would be the use case.”
- P01, Urban, Learn Your Way
@@
The Textbook condition, on the other hand, required more visual focus to process the content in the textbook chapter. Without access to text-to-speech tools or additional support, multitasking within the digital textbook would therefore be challenging. It is worth noting however that there are open questions about the impact of multitasking on the comprehension and recall of information and which tasks are suitable for this (Coens2011). It is therefore crucial to consider approaches to support metacognitive strategies (e.g. reflection) with auditory modalities to ensure effective learning.
8 Discussion
@@
Our results are in support of our hypotheses, which demonstrate that Learn Your Way can lead to both improved learning outcomes and experiences. AI-powered learning platforms can be a powerful driver of pedagogical innovation. In particular, our experiment reveals that these tools are not merely supplementary aids but can play a critical role in reshaping how students interact with and comprehend academic content. While many studies exploring the impact of AI tools on educational outcomes have been mixed (e.g.,(do2025; Sun2024)), our research demonstrates that Learn Your Way can enhance students’ immediate and long-term recall of information and create a more effective learning environment compared to a traditional textbook.
@@
The experimental results suggest that this can be attributed to several factors. First, learning is a complex and multifaceted process and students exhibit diverse preferences in how they process information and acquire knowledge, including visual, kinesthetic, and reading/writing approaches (mccall2024future). Traditional instructional methods often fail to address these unique needs at scale, leading to suboptimal educational experiences (Villegas2024). Instead, engaging with information through multiple representational formats, such as slide decks, audio lessons, or visual mind-maps, facilitates the construction of a more robust and comprehensive mental model of the content (AINSWORTH1999). This is because diverse representations can clarify complex ideas and highlight distinct facets of a concept. Learn Your Way addresses this need by offering students the ability to select their preferred learning modalities, enabling them to explore concepts from a range of engaging and diverse formats. Second, simplified and multimodal presentations can reinforce learning, leading to greater retention when information is seen, heard, and actively engaged with (clark1991dual). By breaking down content into manageable chunks and allowing students to personalize the presentation of the content in dynamic formats based on their learning preferences, we believe students forged better mental connections between different representations, resulting in better comprehension of the content compared to students who only engaged with a digital textbook.
@@
Drawing from established literature on the benefits of adaptive quizzing and personalized feedback (frank2024influence), this study’s quantitative and qualitative findings underscore their critical role in the learning experience. All students who used Learn Your Way experienced quizzing and feedback. Participants’ qualitative feedback further highlighted the value of these capabilities, with many emphasizing how quizzes aided in testing comprehension and retention, while the associated feedback was instrumental in helping them understand the rationale behind their answers.
@@
Given the power of content transformation capabilities on student learning, one might wonder why there was no direct correlation between engagement with usage of more content transformation modalities and better learning outcomes among students who used Learn Your Way. This analysis was exploratory and the pattern of data is in the direction we expected (i.e., more use of Learn Your Way modalities was correlated with better learning outcomes) but not statistically significant. We expect this may have occurred for two reasons. First, this analysis required us to segment our Learn Your Way students into two different groups, which significantly reduced our sample size and therefore power to detect statistical significance. Second, our qualitative feedback suggests that the usage of additional modalities, such as Slides and Video were often treated as supplementary resources for content review, following primary engagement with the Immersive Text and QuizMe features.
@@
“I feel like [Slides] would be good for review after I already read the Immersive Text if I wanted to review the main points.” - P72, Rural, Learn Your Way
@@
Our findings therefore suggest two potential explanations for the observed phenomena. First, a larger sample size could provide a more robust basis for investigating whether increased use of content transformation capabilities positively impacts learning outcomes. Second, it may be the case that a greater quantity of engagement with content transformation features does not, in itself, lead to improved learning outcomes. Rather, our results corroborate previous research demonstrating that enhanced learning is contingent upon the strategic alignment of an individual’s learning preferences with the specific demands of the task (Cordova1996; schneider2018autonomy). In other words, the observed benefits are likely driven by the availability of multimodal options and the resulting learner agency, rather than a cumulative effect of exposure to multiple formats. As our current design did not manipulate modality exposure experimentally, we refrain from implying that using a higher number of transformations leads to superior outcomes. Instead, the results suggest that providing a diverse “menu” of representations allows students to select the single most effective format for their individual needs. While these tools offer significant value, their effectiveness is maximized when learners have the autonomy to select the specific transformations that align with their personal learning goals and contextual needs. Crucially, our data indicates that effectiveness is not predicated on the number of modalities utilized, but rather on the student’s ability to find a representation that aligns with their specific learning needs.
Furthermore, it is important to note that Learn Your Way is not predicated on the efficacy of a single content representation, but rather on the theoretical synergy between learner agency and the benefits of multiple representations (as grounded in Dual Coding Theory and Multimedia Learning principles). Therefore, attempting to isolate the performance contribution of any individual content transformation from the element of student choice would fundamentally detach the intervention from its core theoretical foundation, misrepresenting the mechanism by which we hypothesize the system promotes positive performance outcomes.
@@
Our results also demonstrate that students who engaged with Learn Your Way reported a more positive learning experience compared to those who used a traditional textbook. The present study’s findings are supported by a body of literature suggesting that transforming educational content into multiple representations can make the learning process more enjoyable, engaging, and motivating (AINSWORTH1999). This is particularly powerful when students are able to select representations that best match their individual learning preferences (surjono2015effects; Popescu2010). Learn Your Way, which is capable of dynamically transforming content, represents a critical evolution from standard digital textbooks by enabling more tailored and engaging educational experiences for students.
@@
8.1 Limitations and Future Work
While our findings provide evidence that AI-powered content transformation tools can positively impact learning outcomes and experiences, our study is not without limitations. First, our study took place in an experimental lab setting. Future research should explore best practices for how and when to utilize these content transformation tools within a classroom environment, and collect longitudinal outcomes on learning and retention. Second, a limitation of the current study design is the lack of a feature-matched control group. While the Textbook condition represents the current standard for digital learning, it lacks the interactive scaffolds (e.g., quizzes and feedback) present in Learn Your Way. Consequently, the observed gains may be attributed to the combined effect of content transformation and these interactive elements. Future research should employ a control condition that includes non-AI-generated quizzes to isolate the specific impact of the generative content transformations. Third, while the observed effect sizes for immediate ($r=.24$) and long-term recall ($r=.25$) are characterized as small to medium by traditional conventions, their practical significance in a classroom setting is noteworthy.
In educational contexts, small improvements in recall can lead to significant cumulative gains in student mastery over a full academic year. Furthermore, these effects were achieved with a relatively brief intervention (up to 40 minutes). For a teacher, moving a student from an average score of $8.2/12$ to $9.2/12$ – as seen in our immediate recall results – represents the difference between a “D” and a “C” grade. Given that Learn Your Way is an automated, AI-driven tool, these gains suggest a scalable method for providing personalized support that would otherwise require intensive one-on-one tutoring. Thus, the tool demonstrates a “low-effort, high-yield” potential for augmenting standard classroom instruction.
Finally, while this study provides a foundational proof-of-concept, several factors present potential threats to its external validity. Specifically, the controlled lab environment and the focus on a single age group (high school students) and a single subject area (neuroscience) limit the generalizability of our findings to more ecologically valid settings. To address this, future research should transition from the lab to a longitudinal field study within diverse classroom environments. This will allow researchers to observe how the efficacy of AI-powered learning persists across various disciplines – such as mathematics or literature – where the nature of “transformation” may differ structurally. Furthermore, by expanding the participant pool to include middle school and university students, researchers can identify the specific developmental stages at which these interactive representations provide the most significant pedagogical benefit.
9 Conclusion